# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v3

@@ 3.3.2 When the eighth quadric is not regular
The case (++)(++)( + + ), in which both triples of axes and spears are linearly independent, clearly produces regular matrices, hence the eighth quadric is regular. This section considers what kinds of degenerate quadrics arise in the other cases. We remark that by equation (29) the $\lambda^{0n}_{ij}$ meet in a point exactly when the $x_{n}$ share a common line, i.e., when the resulting matrix $D$ is degenerate. It must, furthermore, have rank $3$ because the $\lambda^{0n}_{ij}$ do not coincide and $o$ is not incident to $O$ and thus is independent of the $x_{n}$. Therefore in case (+)(+-)( + - ), if the $\lambda^{0n}_{ij}$ are dependent, $\mathbf{{T}}^{0}$ is a cone. In case (+)(-+)( - + ), we dualize and obtain a conic. The final point-wise quadric in ring contact is the plane of this conic, doubled. In case ()(--)( - - ), the final point-wise quadric is again a double plane but now it does not arise as the plane of a conic but rather as a contact element together with a pair of lines: the lines of $X$-contact (note however that the $\mathbf{{S}}_{j}$ are not in $X$-contact with each other: as discussed in, e.g. [BO20], there are several different contact relations for which two quadrics may meet in a pair of lines, $X$-contact being the special case where this is the entirety of the intersection; in the present case, this will hold for each $\mathbf{{S}}_{j}$ with the double plane $\mathbf{{T}}^{0}$ but not for the $\mathbf{{S}}_{j}$ with each other). From this it is clear that the full information of the quadrics is not contained in the point-wise or plane-wise description alone but requires knowledge of both (as well as possibly lines). However, in the present paper we leave the precise formulation somewhat one-sided, leaving a more principled exposition to future work.
@@ 4 Proof of Extrusion Lemma
To begin with, the scaling of the matrix representing $\mathbf{{S}}_{0}$ can be chosen arbitrarily and then the matrices representing $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$, and $\mathbf{{T}}^{3}$ are uniquely defined by equation (40). Then it only remains to scale $\mathbf{{S}}_{1},\mathbf{{S}}_{2}$ and $\mathbf{{S}}_{3}$ so that equation (40) holds for both double contacts of each. Let $A,B$ and $C$ be the appropriately scaled matrices representing $\mathbf{{S}}_{0},\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{2}$ respectively:
@@
Because of condition (ii) of the hypothesis of the 8-conic theorem, all the chords of contact for the face $(\mathbf{{S}}_{0},\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{S}}_{3})$ share a point whose representative in the basis $(P_{1},P_{2},P_{3})$ is necessarily in the common kernel of $V^{1}_{3}$, $V^{2}_{0}$ ,$V^{1}_{0}$ and $V^{2}_{3}$. However, by the regularity assumption on $\mathbf{{T}}^{2}$, $C$ has trivial kernel which is only consistent with equation (40) in the case $r=1$.
@@ 4.1 Proof of the 8-quadric theorem with a common contact element
Given a configuration of conics in double contact, equation (39) exhausts all possible extensions of $B$ to a quadric polarizing $O$ and $o$ and having all of these double contacts as slices of ring contacts, with $O$ as the point of concurrency of the contact axes and thus $o$ as the common plane of the ring points. Therefore, to prove the case of the 8-quadric theorem where $o$ is incident to $O$, we apply the 8-conic theorem to a planar slice of the 7-quadric configuration and then extrude the resulting complete 8-conic via equation (39), choosing the parameter $t$ that extrudes the seven conics obtained from the slicing process back to the seven quadrics we started with. The eighth extruded quadric is then the desired conclusion of the 8-quadric theorem. ∎
@@ 5 Examples of 8-quadric configurations
The aim of this section is to illustrate the structure of the 8-quadric theorem by constructing several examples step by step and, by slicing them, to obtain all the famous 2D-porisms mentioned in the introduction. Though the 8-quadric theorem was stated for regular quadrics, much of the reasoning in section 3.3 can be traced through degenerate cases like the ones in this present section. We provide some indications of necessary modifications in the logic of the proof in order to apply to these examples, however, we leave the delineation of the general conditions for allowable degeneracies, and the complete enumeration of the possible cube configurations that may result, to a future work dedicated solely to that subject.
@@
In the next example we consider three spheres $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ such that no sphere is completely contained in another one. All spheres intersect the plane at infinity in the same imaginary circle $\mathbf{{S}}_{0}$, the absolute of Euclidean geometry. Thus each sphere $\mathbf{{T}}^{i}$ is in ring contact with $\mathbf{{S}}_{0}$, where the quadric $\mathbf{{S}}_{0}$ as an imaginary degenerate conic represents at the same time the rings $\mathbf{{R}}_{0}^{i}$. We could proceed as in the previous example by choosing as quadrics $\mathbf{{S}}_{k}$ the intersection of the spheres $\mathbf{{T}}^{i}\cap\mathbf{{S}}_{0}$, which are circles (besides the absolute imaginary circle at infinity) and by slicing with a plane through the centers of the 3 spheres we would obtain again a Brianchon-like theorem, which turns out to be the theorem of the radical center of 3 circles.
Instead we choose cones enveloping each pair of spheres. If the spheres are disjoint there are 6 such cones in total (dual to the 6 conics in Example 5.1) and again there are two cases for the choice of a triplet of cones. In one case the cones share 2 tangent planes which are also common tangent planes of the 3 spheres. See Figure 9, left. Thus the final quadric $\mathbf{{T}}^{0}$ is this pair of planes and the rings $\mathbf{{R}}_{i}^{0}$ are 3 line pairs. This 3D-configuration was already used by G. Monge (according to [DW92]p.154) in order to prove a theorem nowadays bearing his name: The three pairs of external tangents of three circles meet in three points on a line.
In the other case the cones do not admit common tangent planes. See Figure 9, right. Now the rings $\mathbf{{R}}_{i}^{0}$ are obtained by intersecting each cone $\mathbf{{S}}_{i}$ with the corresponding plane $p_{i}^{0}$ which is the join of a pair of vertical axes. The final quadric is again guarantied by lemma 3.4. A planar slice through the centers of the spheres shows that 3 pairs of internal tangents of 3 circles are circumscribed to a conic. This theorem seems to be known only for the case, where the final conic degenerates into a point pair, see [Bak36],[Sal17].
@@ 6 Related research and outlook for future work
We recently became aware of an independent discovery of the 8-conic theorem in [Eid13], Sec. 9.2.
The research leading to this article has also produced two companion articles which provide a different perspective on the material. [CGN] presents a proof of the 8-conic theorem working within the 5D parameters space of conics. [ACG+] does the same for the 8-quadric theorem within the 9D space of quadrics. The proofs have a similar structure to each other, but advantages of the 3D parameter-space approach over the 2D approach become clear in the final stages of the proofs.
This article and [ACG+] are the result of a collaboration with Roger Penrose that arose after the authors of [CGN] contacted him to report on their results. A collaboration ensued that led, among other results, to the successful reconstruction of the original 3D quadric proof of Roger Penrose in the two forms mentioned above.
Roger Penrose also told us that as an undergraduate he first recognized the cube structure of the theorem from examples in which most of the conic sections are degenerate. He then developed a method for gluing together several such cubes to generate new cubes that contain more regular conics. In this way, he finally arrived at the formulation of the general theorem. Furthermore, the possibility of gluing several cubes together is also interesting for obtaining further theorems. For example, this method can be used to extend Poncelet’s theorem for triangles to general polygons. Another example is Neville’s theorem about three ellipses that have a common focal point in pairs: Here, a first application of the eight-conic theorem can be used to show that all three ellipses must have double contact with another conic. This then serves as a starting point for further cubes, attached to the first one. We are currently working on examples of this type of iterated application of the theorem.
Regarding cube configurations with degenerate conics or quadrics, further conditions are needed to ensure the validity of the main theorems, presented in this article. This can be investigated in parameter space by looking at limits of regular configurations. It seems very worthwhile to us to further develop these techniques making use of the projective geometric structures of parameter space, as done in [CGN] and [ACG+].

### v4

@@ 3.4.2 When the eighth quadric is not regular
The case (++)(++)( + + ), in which both triples of axes and spears are linearly independent, clearly produces regular matrices, hence the eighth quadric is regular. This section considers what kinds of degenerate quadrics arise in the other cases. We note that by Eq. (29) the $\lambda^{0n}_{ij}$ meet at a point exactly when the $x_{n}$ share a common line, i.e., when the resulting matrix $D$ is degenerate. It must also have rank $3$ because the $\lambda^{0n}_{ij}$ do not coincide and $o$ is not incident to $O$ and thus is independent of the $x_{n}$. Therefore, in case (+)(+-)( + - ), if the $\lambda^{0n}_{ij}$ are dependent, $\mathbf{{T}}^{0}$ is a cone. In case (+)(-+)( - + ), we dualize and obtain a conic. The final point-wise quadric in ring contact is the plane of this conic, doubled. In case ()(--)( - - ), the final point-wise quadric is again a double plane but now it does not arise as the plane of a conic but rather as a contact element together with a pair of lines: the lines of $X$-contact (note however that the $\mathbf{{S}}_{j}$ are not in $X$-contact with each other: as discussed in, e.g. [BO20], there are several different contact relations for which two quadrics may meet in a pair of lines, $X$-contact being the special case where this is the entirety of the intersection; in the present case, this will hold for each $\mathbf{{S}}_{j}$ with the double plane $\mathbf{{T}}^{0}$ but not for the $\mathbf{{S}}_{j}$ with each other). From this it is clear that the full information of the quadrics is not contained in the point-wise or plane-wise description alone but requires knowledge of both (as well as possibly lines). However, in the present paper we leave the precise formulation somewhat one-sided, leaving a more principled exposition to future work.
@@ 4 Proof of Extrusion Lemma
To begin with, the scaling of the matrix representing $\mathbf{{S}}_{0}$ can be chosen arbitrarily and then the matrices representing $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$, and $\mathbf{{T}}^{3}$ are uniquely defined by Eq. (40). Then it only remains to scale $\mathbf{{S}}_{1},\mathbf{{S}}_{2}$ and $\mathbf{{S}}_{3}$ so that Eq. (40) holds for both double contacts of each. Let $A,B$ and $C$ be the appropriately scaled matrices representing $\mathbf{{S}}_{0},\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{2}$ respectively:
@@
Because of condition (ii) of the hypothesis of the 8-conic theorem, all the chords of contact for the face $(\mathbf{{S}}_{0},\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{S}}_{3})$ share a point whose representative in the basis $(P_{1},P_{2},P_{3})$ is necessarily in the common kernel of $V^{1}_{3}$, $V^{2}_{0}$ ,$V^{1}_{0}$ and $V^{2}_{3}$. However, by the regularity assumption on $\mathbf{{T}}^{2}$, $C$ has trivial kernel which is only consistent with Eq. (40) in the case $r=1$.
@@ 4.1 Proof of the 8-quadric theorem with a common contact element
Given a configuration of conics in double contact, Eq. (39) exhausts all possible extensions of $B$ to a quadric polarizing $O$ and $o$ and having all these double contacts as slices of ring contacts, with $O$ as the point of concurrency of the contact axes and thus $o$ as the common plane of the ring points. Therefore, to prove the case of the 8-quadric theorem where $o$ is incident to $O$, we apply the 8-conic theorem to a planar slice of the 7-quadric configuration and then extrude the resulting complete 8-conic via Eq. (39), choosing the parameter $t$ that extrudes the seven conics obtained from the slicing process back to the seven quadrics with which we started. The eighth extruded quadric is then the desired conclusion of the 8-quadric theorem. ∎
@@ 5 Examples of 8-quadric configurations
The aim of this section is to illustrate the structure of the 8-quadric theorem by constructing several examples step by step and, by slicing them, to obtain all the famous 2D-porisms mentioned in the Introduction. Although the 8-quadric theorem was stated for regular quadrics, much of the reasoning in Sect. 3.4 can be traced through degenerate cases such as those in this section. We provide some indications of necessary modifications in the logic of the proof in order to apply to these examples; however, we leave the delineation of the general conditions for allowable degeneracies, and the complete enumeration of the possible cube configurations that may result, to a future work dedicated solely to that subject.
@@
In the next example, we consider three spheres $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ such that no sphere is completely contained in another. All spheres intersect the plane at infinity in the same imaginary circle $\mathbf{{S}}_{0}$, the absolute of Euclidean geometry. Thus, each sphere $\mathbf{{T}}^{i}$ is in ring contact with $\mathbf{{S}}_{0}$, where the quadric $\mathbf{{S}}_{0}$ as an imaginary degenerate conic represents at the same time the rings $\mathbf{{R}}_{0}^{i}$. We could proceed as in the previous example by choosing as quadrics $\mathbf{{S}}_{k}$ the intersection of the spheres $\mathbf{{T}}^{i}\cap\mathbf{{S}}_{0}$, which are circles (besides the absolute imaginary circle at infinity) and by slicing with a plane through the centers of the 3 spheres we would obtain again a Brianchon-like theorem, which turns out to be the theorem of the radical center of 3 circles.
Instead we choose cones enveloping each pair of spheres. If the spheres are disjoint there are 6 such cones in total (dual to the 6 conics in Example 5.1) and again there are two cases for the choice of a triplet of cones. In one case the cones share 2 tangent planes, which are also common tangent planes of the 3 spheres. See Figure 9, left. Thus the final quadric $\mathbf{{T}}^{0}$ is this pair of planes and the rings $\mathbf{{R}}_{i}^{0}$ are 3 line pairs. This 3D configuration was already used by G. Monge (according to [DW92]p.154) in order to prove a theorem today bearing his name: The three pairs of external tangents of three circles meet in three points on a line.
In the other case the cones do not admit common tangent planes. See Figure 9, right. Now the rings $\mathbf{{R}}_{i}^{0}$ are obtained by intersecting each cone $\mathbf{{S}}_{i}$ with the corresponding plane $p_{i}^{0}$ which is the join of a pair of vertical axes. The final quadric is again guaranteed by Lemma 3.4. A planar slice through the centers of the spheres shows that 3 pairs of internal tangents of 3 circles are circumscribed to a conic. This theorem seems to be known only for the case, where the final conic degenerates into a point pair, see [Bak36],[Sal17].
@@ 6 Related research and outlook for future work
We recently became aware of an independent discovery of the 8-conic theorem in [Eid13], Sect. 9.2.
The research leading to this article has also produced two companion articles which provide a different perspective on the material. [CGN] presents a proof of the 8-conic theorem working within the 5D parameters space of conics. [ACG+] does the same for the 8-quadric theorem within the 9D space of quadrics. The proofs have a similar structure to each other, but the advantages of the 3D parameter space approach over the 2D approach become clear in the final stages of the proofs.
This article and [ACG+] are the result of a collaboration with Roger Penrose that emerged after the authors of [CGN] contacted him to report on their results. A collaboration ensued that led, among other results, to the successful reconstruction of the original 3D quadric proof of Roger Penrose in the two forms mentioned above.
The final author recalls that as an undergraduate he first recognized the cube structure of the theorem from examples in which most of the conic sections are degenerate. He then developed a method for gluing together several such cubes to generate new cubes that contain more regular conics. In this way, he finally arrived at the formulation of the general theorem. Furthermore, the possibility of gluing several cubes together is also interesting to obtain further theorems. For example, this method can be used to extend Poncelet’s theorem for triangles to general polygons. Another example is Neville’s theorem about three ellipses that have a common focal point in pairs: Here, a first application of the eight-conic theorem can be used to show that all three ellipses must have double contact with another conic. This then serves as a starting point for further cubes, attached to the first one. We are currently working on examples of this type of iterated application of the theorem.
Regarding cube configurations with degenerate conics or quadrics, further conditions are needed to ensure the validity of the main theorems presented in this article. This can be investigated in parameter space by looking at limits of regular configurations. It seems very worthwhile to us to further develop these techniques making use of the projective geometric structures of parameter space, as done in [CGN] and [ACG+].